Find algebraic expression. The perimeter of a rectanle, given that the length is x yards and the width is 10 yards shorter
Answer by stanbon(75887)

Find algebraic expression. The perimeter of a rectanle, given that the length is x yards and the width is 10 yards shorter
---------------
Form:
P = 2(L + W)
P = 2(x + x-10)
p = 2(2x-10)
p = 4x-20
=============
Cheers,
Stan H.